Find the x and y Intercepts of the eqution 3x-2x=9
and Find the slope and y intrcept of the grapgh of 2x+3y=-4

Answer by queenofit(120)

I am not sure you put the first equation correct since there are 2 x values.
but if you use
3x-2y=9 I can do that one
first put it in y=mx+b form
-2y=-3x+9
y=3/2x+9 your slope is m or 3/2 in this case
to find the x intercept set y = 0 and solve for x
to find the y intercept set x=0 and solve for y
0=3x+9
-9=3x
-3=x (-3,0) is x intercept
_____________________________________
-2y=0+9
-2y=9
y=-9/2 (0,-9/2) is y intercept

____________________________________________________________________________
find slope and y intercept of 2x+3y=-4
first put in y=mx+b form. your slope is always the m value
3y=2x-4
y=2/3x-4 slope is 2/3
now set x=0 and solve for y= yintercept
you can do it use my example above......